-3d -1.2 =0.9
What is d?

Answer by tommyt3rd(5050)

-3d -1.2 =0.9
-3d-1.2+1.2=0.9+1.2
-3d=2.1
-3d/(-3)=2.1/(-3)
d=-0.7